Which options that earns a higher interest?
Option 1: a person invested $4000, he is offered an interest of 5,7 (compunded monthly) for two years.

Option 2: a person invested $4000, he is offered an interest of 5,9(compunded annually) for two years.
1
Expert's answer
2020-02-21T10:15:39-0500

Option 1: "A=4000(1+\frac{0.057}{12})^{2*12}=4481.80."

Option 2: "A=4000(1+0.059)^2=4485.92."

Option 2 earns a higher interest.